OpenStudy (anonymous):

Find three solutions of the equation. y = –2x a.(0, 0), (3, –8), (–2, –4) b.(1, –2), (0, 0), (–1, 1) c.(–2, 4), (1, –2), (0, 0) d.(–2, 4), (1, –2), (2, –3)

OpenStudy (anonymous):

A tip. Substitute the x value and the y value and see if LHS = RHS. Eg In y =-2x Substitute 1) (0,0) 0 = -2(0) 0 = 0 LHS = RHS Hence correct that (0,0) lies on the line Eg 2 (3,-8) -8 = -2(3) -8 = -6 ===> False So (3,-8) is not a point on the line y=-2x So option A is invalid Similarly try for options B, C, D

OpenStudy (anonymous):

just plug those points in each choice. if they all result in a true statement, then that's the choice.

OpenStudy (anonymous):

Find three solutions of the equation. Y=-2x-1